OpenStudy (codym251):

Drag the fractions to put them in order from least to greatest. 1/4 1/5 1/8 1/10

OpenStudy (zeus4263):

flip the entire list

OpenStudy (zeus4263):

bottom is smallest top is largest

OpenStudy (zeus4263):

in that order

OpenStudy (qwertty123):

LEAST TO GREASTEST -1/10 -1/8 -1/5 -1/4

OpenStudy (zeus4263):

1/10 - Smallest 1/8 1/5 1/4 - Largest

OpenStudy (anonymous):

tr turning fractions to decimals

OpenStudy (zeus4263):

so as i said, flip the list

OpenStudy (anonymous):

typo...try

OpenStudy (qwertty123):

the higher the denominator the lower the number is

OpenStudy (zeus4263):

exactly

OpenStudy (zeus4263):

so, Flip. The. List.

OpenStudy (codym251):

thanks guys

OpenStudy (codym251):

here other one. 2/3 1/4 1/6 7/8

OpenStudy (qwertty123):

LEAST TO GREASTEST 1/6 1/4 2/3 7/8

OpenStudy (qwertty123):

please only post 1 question per post- Thanks QUACK (^)<

OpenStudy (codym251):

i got 5 more to do. can you help me?

OpenStudy (qwertty123):

and if you don't mind medaling me since i helped with both and sure just ask each one as a different post

OpenStudy (codym251):

1/2 3/4 1/5 7/10

OpenStudy (codym251):

hello??????????????

OpenStudy (xxdragoonxx):

1/5 = 0.2 1/2 = 0.5 7/10 = 0.7 3/4 = 0.75

OpenStudy (codym251):

it's not decimals

OpenStudy (xxdragoonxx):

It's how you can order the fractions from least to greatest that way also.

OpenStudy (codym251):

oh